A gas has a density of 1.73 g L-1 at 52 °C and P = 1.06 atm. What will be its density at 169 °C? Calculate the molar mass of the gas, assuming it obeys the ideal gas law.

Density at 169 °C = ______ g L-1

Molar mass = _______g mol-1
Answer #1

Step 1: calculate molar mass

P = 1.06atm

T= 52.0 oC

= (52.0+273) K

= 325 K

density = 1.73 g/L

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

1.06 atm * MM = 1.73g/L * 0.08206 atm.L/mol.K *325.0 K

MM = 43.53 g/mol

Step 2:Now calculate density

P = 1.06atm

T= 169.0 oC

= (169.0+273) K

= 442 K

Use:

p*molar mass=density*R*T

Put Values:

1.06 atm *43.53 g/mol = density * 0.08206 atm.L/mol.K *442.0 K

density = 1.27 g/L

Answer:

1) 1.27 g/L

2) 43.53 g/mol